OpenStudy (anonymous):

Find the real or imaginary solutions of each equation by factoring. 125x^3-27=0

OpenStudy (anonymous):

okay i have it but its long k

OpenStudy (anonymous):



OpenStudy (anonymous):

First notice: \[125x^3-27=(5x)^3-3^3\] Difference of cubes: \[a^3-b^3=(a-b)(a^2+ab+b^2)\] Set \(a=5x\) and \(b=3\).